Hypotenuse of isosceles triangle is five centimeters how long are its legs?
I'm not sure how to go about this...I was told that you couldn't find the length of legs if you only had one measurement.

Answer by scott8148(6628)

what you were told is true ... for "regular" triangles ___ this is a "special" case
___ isosceles means that the two legs are equal

by Pythagoras ___ x^2 + x^2 = 5^2

2x^2 = 25 ___ x^2 = 25/2 ___ x = (5/2) sqrt(2)